OpenStudy (anonymous):

If I have a 1x1x1 cube. How can I make it so it has twice the surface area?

OpenStudy (anonymous):

I think 6

OpenStudy (solomonzelman):

well, the surface Area of a cube the dimensions of which are 1, 1, 1 is ?

OpenStudy (solomonzelman):

There are 6 same faces, 1 by 1 each. So if each face has area of 1 (meter squared), then all of these 6 faces added would give you the (entire surface) area of ?

OpenStudy (anonymous):

6

OpenStudy (solomonzelman):

yes, the surface area of a cube whose side (each one) is 1, is 6.

OpenStudy (anonymous):

so how do you make it so it is twice that?

OpenStudy (solomonzelman):

Now, how did we obtain this surface area (once again) ? (I will denote my side as "s" ) \(\large\color{black}{ \displaystyle {\rm A_{~surface~of~a~cube}}~=~6s^2 }\) do you understand this formula?

OpenStudy (solomonzelman):

(each face is s^2, and there are 6 of these faces )

OpenStudy (solomonzelman):

yes or no ?

OpenStudy (anonymous):

yes, so to make it so the surface area is twice as much you square it?

OpenStudy (solomonzelman):

so you have the following for your surface area: \(\large\color{black}{ \displaystyle {\rm A_{~surface~of~a~cube}}~=~6(1)^2 }\) which gave you that A=6

OpenStudy (solomonzelman):

but, we need to go backwards. We want to know how long should each side should be to get a surface area of 12.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm 12}~=~6s^2 }\)

OpenStudy (solomonzelman):

Solve for s (and disregard any negative solution if you get it)

OpenStudy (anonymous):

so 12/6=2 so \[\sqrt{2}=s ^{2}\]

OpenStudy (solomonzelman):

you mean \(\large\color{black}{ \displaystyle s=\sqrt{2} }\) ?

OpenStudy (anonymous):

yea sorry

OpenStudy (solomonzelman):

it is alright, as long as you get it (which you apparently do).

OpenStudy (solomonzelman):

So, to have a surface area of 12 (which is twice more than the surface of 6), you would need each side to be \(\large\color{black}{ \displaystyle \sqrt{2} }\) units long

OpenStudy (solomonzelman):

any questions about this ?

OpenStudy (anonymous):

no, thank you for your help

OpenStudy (solomonzelman):

yw